# Comparing Styles across Languages

### v2

@@ 4 Comparing Politeness via PoliteLex
English |
PoliteLex |
98.0% |
Chinese |
PoliteLex |
83.7% |
Spanish |
MT |
94.4% |
@@ 5 Comparing Politeness via Dialogue Acts
Conventional Opening |
“Hi all, this article is in desperate need of attention.” |
Declarative Statement |
“As it stands, we do not know the place or date of Shapiro’s birth.” |
Opinion Statement |
“There is no need to be combative and attack over typos.” |
Forward-Function: Thanking |
“Looks good, thanks for the help!” |
Yes/No Question |
“Is the second picture an actual representation of the BCG vaccine?” |
Conventional Closing |
“Happy to discuss anything further.” |
@@ Results.
“最后应用一节,有没有必要加入那么多图片?” (“In the last application section, is it necessary to add so many pictures?”)
@@ 6 Ablation Analysis
To investigate the role of LM size and architecture, we fine-tune Llama-2-7b Touvron et al. (2023) to analyze politeness. Comparing the results from Llama-2-7b to the results from XLM-RoBERTa, we notice stability in the direction of importance score (i.e. positive, negative, and neutral lexical categories/dialog acts are stable across both LMs). Interestingly, we observe differences in the magnitude of importance score (e.g. Llama-2-7b sees the ”Greeting” category in Chinese to be a larger indicator of politeness than XLM-RoBERTa does).
@@ A.1 Data Collection
We scrape Wikipedia Editor Talk Pages in English, Spanish, Japanese, and Chinese using Wikimedia555https://dumps.wikimedia.org/backup-index.html and keep the Talk Pages discussing articles that are present in all four languages. Next, we extract utterances, or two–three sentence snippets from a single user, such that the length of a single utterance does not exceed 512 tokens. We anonymize utterances to remove names, replacing mentions with “@user” and hyperlinks, replacing hyperlinks with “¡url¿”. We then randomly sample 5,700 utterances from each language.
@@ A.3 Data Processing
English |
.667 |
.662 |
Spanish |
.738 |
.642 |
Japanese |
.659 |
.650 |
Chinese |
.612 |
.646 |
@@ Appendix C Ablation Analysis: Additional Details
Gratitude |
thanks, thank you, i appreciate |
Deference |
good, great, interesting |
Greeting |
hello, hi, hey |
Apologetic |
sorry, apologize, apologies |
Please |
please, pls, plse |
Please Start |
please |
Indirect (btw) |
by the way, btw |
Direct Question |
what, when, how |
Direct Start |
but, and, so |
Subjunctive |
could you, would you |
Indicative |
can you, will you |
1st Person Start |
i, my, mine |
1st Person Plural |
we, us, our |
1st Person |
i, my, me |
2nd Person |
you, your, u |
2nd Person Start |
you, your |
Hedges |
should, think, seems |
Factuality |
actually, really, in fact |
Emergency |
right now, immediately, at once |
Ingroup Identity |
mate, homie, dude |
Praise |
great, excellent, super |
Promise |
sure, must, certainly |
Together |
together |
Direct ”You” |
you, u |
Positive |
like, well, good |
Negative |
issue, wrong, problem |

### v3

@@ 4 Comparing Politeness via PoliteLex
| English | PoliteLex | 98.0% |
| Chinese | PoliteLex | 83.7% |
| Spanish | MT | 94.4% |
@@ 5 Comparing Politeness via Dialogue Acts
| Conventional Opening | “Hi all, this article is in desperate need of attention.” |
| Declarative Statement | “As it stands, we do not know the place or date of Shapiro’s birth.” |
| Opinion Statement | “There is no need to be combative and attack over typos.” |
| Forward-Function: Thanking | “Looks good, thanks for the help!” |
| Yes/No Question | “Is the second picture an actual representation of the BCG vaccine?” |
| Conventional Closing | “Happy to discuss anything further.” |
@@ Results.
UTF8gbsn“最后应用一节,有没有必要加入那么多图片?” (“In the last application section, is it necessary to add so many pictures?”)
@@ 6 Ablation Analysis
To investigate the role of LM size and architecture, we fine-tune Llama-2-7b Touvron et al. (2023) to analyze politeness. Comparing the results from Llama-2-7b to the results from XLM-RoBERTa, we notice stability in the direction of importance score (i.e. positive, negative, and neutral lexical categories/dialog acts are stable across both LMs). Interestingly, we observe differences in the magnitude of importance score (e.g. Llama-2-7b sees the "Greeting" category in Chinese to be a larger indicator of politeness than XLM-RoBERTa does).
@@ A.1 Data Collection
We scrape Wikipedia Editor Talk Pages in English, Spanish, Japanese, and Chinese using Wikimedia555https://dumps.wikimedia.org/backup-index.html and keep the Talk Pages discussing articles that are present in all four languages. Next, we extract utterances, or two–three sentence snippets from a single user, such that the length of a single utterance does not exceed 512 tokens. We anonymize utterances to remove names, replacing mentions with “@user” and hyperlinks, replacing hyperlinks with “<url>”. We then randomly sample 5,700 utterances from each language.
@@ A.3 Data Processing
| English | .667 | .662 | |
| Spanish | .738 | .642 | |
| Japanese | .659 | .650 | |
| Chinese | .612 | .646 |
@@ Appendix C Ablation Analysis: Additional Details
| Gratitude | thanks, thank you, i appreciate |
| Deference | good, great, interesting |
| Greeting | hello, hi, hey |
| Apologetic | sorry, apologize, apologies |
| Please | please, pls, plse |
| Please Start | please |
| Indirect (btw) | by the way, btw |
| Direct Question | what, when, how |
| Direct Start | but, and, so |
| Subjunctive | could you, would you |
| Indicative | can you, will you |
| 1st Person Start | i, my, mine |
| 1st Person Plural | we, us, our |
| 1st Person | i, my, me |
| 2nd Person | you, your, u |
| 2nd Person Start | you, your |
| Hedges | should, think, seems |
| Factuality | actually, really, in fact |
| Emergency | right now, immediately, at once |
| Ingroup Identity | mate, homie, dude |
| Praise | great, excellent, super |
| Promise | sure, must, certainly |
| Together | together |
| Direct "You" | you, u |
| Positive | like, well, good |
| Negative | issue, wrong, problem |